OpenStudy (anonymous):

multiply (6+ sqrt 5)(4-2 sqrt 5) 24-2 sqrt 5 14-8 sqrt 5 24-10 sqrt 5 14-16 sqrt 5

OpenStudy (anonymous):

14-8sqrt5

OpenStudy (angela210793):

=24-12sqr5+4sqr5-10=14-8sqr5

OpenStudy (anonymous):

\[(6+\sqrt{5)}(4-2\sqrt{5)}\] \[24-12\sqrt{5}+4\sqrt{5}-10\] \[14-8\sqrt{5}\]